1 + 4/7- 3/4
I am not sure where the order of operations start? Do I start with the fractoions and find a common denominator and then what do I do with the 1?

Answer by Theo(13342)

your expression is:
1 + 4/7- 3/4
steps:
1. divide 4 by 7.
2. add it to 1.
3. divide 3 by 4.
4. subtract it from the result in step 2.
the results from your calculator will show:
1. 4/7 = .5714295871
2. .5714295871 + 1 = 1.5714295871
3. 3/4 = .75
4. 1.5714295871 - .75 = .821428571
if you want to deal with fractions rather than decimals, you would need to get everything under a common denominator before performing any operations.
your expression is:
1 + 4/7- 3/4
lowest common denominator appears to be 28, so makes everything equivalent to a fraction with a denominator of 28.
you get:
28/28 + 16/28 - 21/28
since the denominators are common, you can add the numerators together under the common denominator to get:
(28 + 16 - 21) / 28
perform the operation within each set of parenthese first to get:
23/28
decimal equivalent of this is equal to .821428571.

that's the same answer you got before.
here's a reference on order of arithmetic operations.
http://math.about.com/library/weekly/aa040502a.htm